Easy One Part! Will give thumbs up!

2. A balanced subset of data from a completely randomized experiment on germination of barley is summarized in the followingml H,0 1 911. = 10.0 721. = 5.5 Age of Seeds (weeks) 3 6 9 12 712. = 11.5 113. = 22.0 414. = 31.5 715. = 28.5 722. = 2.0 723.(b) (2 pts] Compute 2 5 2 2 5 2 ΣΣΣωμ. - - ΣΣΣνή. - ισ.. 1=1 j=1 k=1 1=1 =1k=1 (where n is the total number of observations
Answer #1

2 5 2 2 do sr Yeny = 6811 – 9x(14. O D = 6811 – 1958,0625 = 4852.9375 lurralde Anno Doni con - 2 2 2 2 5 2 12 => [ 5 2 El Juk2 5 2 2 do sr Yeny = 6811 – 9x(14. O D = 6811 – 1958,0625 = 4852.9375 lurralde Anno Doni con - 2 2 2 2 5 2 12 => [ 5 2 El Juk